OpenStudy (anonymous):

A clothing designer enters a competition to show ten original outfits at a fashion show. If the designer has 12 outfits to choose from, how many ways are there to arrange the order in which he will show ten of them?

OpenStudy (anonymous):

@Brostep0s

OpenStudy (anonymous):

12P10=12/2=12*11=132 variations.

OpenStudy (anonymous):

There are only like 2 more left if you could help me on those last 2 ones!

OpenStudy (anonymous):

post

OpenStudy (anonymous):

or this 12!/(12-2)!=132 @Brostep0s @BradSuga right?

OpenStudy (anonymous):

yes, @megz08 that is the correct way to do it, I am just using simpler terms.,

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

What would that represent in one single sentence?

OpenStudy (anonymous):

what do you mean by represent?

OpenStudy (anonymous):

example: n things can be arranged in n! ways????? like that?

OpenStudy (anonymous):

"write a sentence to explain the meaning of the answer for the situation. "

OpenStudy (anonymous):

do it like @megz08 stated. Just reinstate that pluging in your answers.

OpenStudy (anonymous):

Alright!

OpenStudy (anonymous):

what did you right on the sentence? just checking.